best buy recently charged $17.99 per box of pens and $7.49 per box for pencils.
If Hag High School purchased 120 such boxes for a total of $1234.80 how many boxes of each type did they purchase?

Answer by ewatrrr(24785)

 
Hi,
purchased 120 such boxes
Let x and (120 -x) represent the number of boxes of pens and pencils respectively
Question states***
17.99x + 7.49(120-x) = 1234.80
10.50x = 1234.80 - 7.49*120
10.5x = 336
x = 32 boxes of pens and 88 boxes of pencils (120-32)
CHECKING our Answer***
32*$17.99 + 88*$7.49 = $575.68 + $659.12 = $1234.80